# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

### v1

@@ 4 Challenge Design
The TUS-REC2024 Challenge [^1] https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/ [^2] https://doi.org/10.5281/zenodo.10991500 is designed following the BIAS [36] Reporting Guideline for enhanced quality and transparency of biomedical research. This Challenge is associated with 5th International Workshop of Advances in Simplifying Medical UltraSound (ASMUS) at MICCAI 2024. The training and validation datasets are publicly available under CC BY-NC-SA license. The Challenge is an open-ended Challenge, and submissions are welcome even after the official deadline. The test set remains held out and will be used exclusively for benchmarking reconstruction performance.
@@ 4.1 Task Description
Aiming at estimating the location for each ultrasound frame in 3D space, this Challenge is tasked to predict four different sets of transformation-representing dense displacement field (DDF), a set of displacement vectors on individual pixels and a set of displacement vectors on provided landmarks, at both global and local levels. The global level DDFs denote the displacement between the current frame and the first frame, and the local level DDFs represent the displacement between the current frame and the previous frame. There are no restrictions on the internal design of the algorithm, for example, whether it is learning-based; processes data at the frame, sequence, or scan level; or assumes rigid, affine, or non-rigid transformations.
Participating teams are provided with sequential data and may choose to leverage its spatiotemporal information if useful. Each team’s model should take an ultrasound scan as input and output four sets of pixel-wise displacement vectors, representing the transformations to a reference frame (i.e., the first frame or the previous frame in the sequence). During evaluation, the submitted dockerized models will be used to generate these displacement fields, from which accuracy scores will be computed to assess reconstruction performance at both local and global levels.
@@ 4.2.1 Data Collection
The dataset [^3] https://doi.org/10.5281/zenodo.11178508444https://doi.org/10.5281/zenodo.11180794555https://doi.org/10.5281/zenodo.11355499666https://doi.org/10.5281/zenodo.12752245 used in this Challenge was collected from both the left and right forearms of 85 volunteers at University College London (UCL), United Kingdom. This study was performed in accordance with the ethical standards in the 1964 Declaration of Helsinki and its later amendments or comparable ethical standards. Approval was granted by the Ethics Committee of local institution on 20th Jan. 2023 [24055/001]. The subject cohort was diverse in terms of race, gender, and age. Fig. 2 illustrates the equipment setup used during data acquisition. There were no specific exclusion criteria, except for individuals with allergies or skin conditions that could be aggravated by the ultrasound gel. All scanned forearms were confirmed to be in healthy condition.
@@
2D ultrasound images were acquired using an Ultrasonix machine (BK, Europe) equipped with a curvilinear probe (4DC7-3/40). The ultrasound frames were captured at a rate of $20$ frames per second, with a resolution of $480\times 640$ pixels, without speckle reduction. Imaging was performed at a frequency of $6$ MHz, with a dynamic range of $83$ dB, an overall gain of $48\%$, and a depth of $9$ cm. Both the left and right forearms of volunteers were scanned. For each forearm, the ultrasound probe was moved along three trajectories, Straight line shape, C shape, and S shape, in both distal-to-proximal and proximal-to-distal directions. These scans were performed with the ultrasound plane perpendicular of and parallel to the scanning direction.
The selection of an appropriate tracking system is determined by the specific clinical environment and the required level of spatial accuracy. In this study, an optical tracking system was chosen due to its greater accuracy and operational convenience compared to EM tracking systems. Specifically, the NDI Polaris Vicra (Northern Digital Inc., Canada) was employed. During acquisition, position data recorded by the optical tracker was captured by the PLUS Toolkit [30], alongside the ultrasound images. For each ultrasound frame, the system records the pose of the tracker tool in the camera coordinate system, represented as a homogeneous transformation matrix. The calibration matrix was obtained using a pinhead-based method [37], defining the transformation between the ultrasound image coordinate system and the tracker tool coordinate system at the time of data acquisition. The data is temporally calibrated, with timestamps aligned between the transformations recorded by the optical tracker and the ultrasound frames acquired from the ultrasound machine, as described previously in Section 3.2.
@@ 4.2.2 Sources of Errors
The primary source of error arises from the precision limitations of the optical tracker. All labels were obtained using an optical tracker with the 3D root-mean-square (RMS) volumetric accuracy acceptance criterion being less than or equal to 0.25 mm and the 3D RMS repeatability acceptance criterion being less than or equal to 0.20 mm. Slight forearm movements may occur during scanning, which is also expected in clinical environments where this technique would be deployed. These motion-induced errors are assumed to be random across different cases. Their impact will be accounted for in the statistical analysis during result summarisation and interpretation. Additional sources of error include inaccuracies in the calibration process (both spatial and temporal), pressure-induced skin deformation, as well as the intra- and inter-observer variability during ultrasound data acquisition, affecting probe positioning and image capture quality. These levels of error are much lower than the typical state-of-the-art reconstruction errors in this application, which has been widely reported to range from several millimeters to tens of millimeters.
@@ 4.2.3 Data Pre-processing
For each scan, ultrasound frames with invalid transformation matrices, which were typically caused by blocked line of sight, were excluded. The remaining raw images, along with their corresponding transformation matrices, were temporally ordered and stored as key-value records in a .h5 file.
@@ 4.2.4 Data Split
Statistical power analysis was performed to determine the appropriate test sample size and minimise the likelihood of Type I and Type II errors in hypothesis testing. The effect size was calculated using Cohen’s D value, where the system error of the optical tracker ($0.25$ mm) was considered the meaningful difference between group means, and the standard deviation ($0.46$ mm) was derived from the results reported by [38]. A statistical power analysis for a t-test, assuming a significance level of $0.05$ and a statistical power of $0.9$, indicated a required test sample size of $31$. To ensure adequate power, we rounded up to $32$ samples ($768$ scans in total). This setup limits the probability of a Type II error to $10\%$ and a Type I error to $5\%$.
The dataset was randomly split into training, validation, and test sets, comprising $50$, $3$, and $32$ subjects, respectively. This corresponds to $1200$, $72$, and $768$ scans, totaling 606597, 34746, and 384105 frames. Ultrasound scans from the same subject will be assigned to the same set which avoids the information leak. Detailed information is described in Table 2. Specifically, the structure of the validation dataset is the same as that of the test set to ensure compatibility with the pre-defined folder hierarchy and naming conventions. This design allows submitted Docker images to run seamlessly on the test set and also supports their use for parameter tuning during validation.
@@
Train Validation Test
Subjects 50 3 32
Scans 1200 72 768
Frames 606597 34746 384105
Straight line shape subjects 50 3 32
C shape subjects 50 3 32
S shape subjects 50 3 32
Straight line shape scans 400 24 256
C shape scans 400 24 256
S shape scans 400 24 256
Straight line shape frames 192117 10515 119421
C shape frames 202654 11655 128721
S shape frames 211826 12576 135963
Parallel scanning subjects 50 3 32
Perpendicular scanning subjects 50 3 32
Parallel scanning scans 600 36 384
Perpendicular scanning scans 600 36 384
Parallel scanning frames 298722 17228 188399
Perpendicular scanning frames 307875 17518 195706
Left arm subjects 50 3 32
Right arm subjects 50 3 32
Left arm scans 600 36 384
Right arm scans 600 36 384
Left arm frames 301155 17081 192118
Right arm frames 305442 17665 191987
Distal-to-proximal scanning subjects 50 3 32
Proximal-to-distal scanning subjects 50 3 32
Distal-to-proximal scanning scans 600 36 384
Proximal-to-distal scanning scans 600 36 384
Distal-to-proximal scanning frames 298803 16908 181844
Proximal-to-distal scanning frames 307794 17838 202261
@@ 4.3.1 Metrics Definition
We use DDFs to evaluate the reconstruction performance, borrowing the widely recognised term used in non-rigid image registration for clarity and intuition. For each scan, participating methods are tasked to generate two types of displacement vectors representing frame-to-frame transformations, hereinafter referred to as predictions, at both global and local levels: 1) global displacement vectors are used to reconstruct all frames (excluding the first) relative to the first frame of the scan, which serves as the global reference frame; 2) local displacement vectors are used to reconstruct each frame (excluding the first) relative to its immediately previous frame, which serves as the local reference frame.
The performance of each submitted method will be assessed for every scan using two metrics: landmark reconstruction error and pixel reconstruction error: 1) landmark reconstruction error is defined as the average Euclidean distance between the ground-truth-reconstructed frame and the prediction-reconstructed frame, computed over a predefined set of landmarks. 2) pixel reconstruction error is similarly defined as the average Euclidean distance between the ground-truth and predicted reconstructions, calculated over all pixels in every frame except the first.
Accordingly, each method should produce the following four sets of displacement vectors:
@@
Specifically, runtime will be included as an additional evaluation metric. It is defined as the consumed time of predicting the positions for all frames but the first frame in a scan, averaged across all scans in the test set. The scale-invariant feature transform (SIFT) [39] algorithm was applied to detect landmarks. For each scan, 20 landmarks with the highest response values were selected.
@@ 4.3.2 Rationale of Evaluation Metrics
Use of Euclidean distance-based error metrics vs. transformation parameter-based errors. Direct evaluating the accuracy of parameters of transformation matrix presents challenges, as the relative influence and weighting of rotational and translational components can vary significantly depending on experimental setups, imaging configurations, reference coordinate systems, and definitions of rotational axes. These factors are also often application-dependent. Therefore, this Challenge adopts Euclidean distance-based metrics, which offer a more direct and practical assessment of the discrepancy between ground truth and predicted positions in physical space.
Use of displacement-based transformation representations vs. rigid / affine matrices. Although ground-truth transformations are provided in the form of rigid transformation, we argue, based on practical experience in developing similar numerical algorithms, that requiring submissions to output homogeneous transformation matrices is not only unnecessary, but sometimes misleadingly encourages a more numerically challenging solution due to issues such as gimbal lock in using rotation matrix, local minima in numerical optimisation. In contrast, displacement-based representations allows flexibility for a quantitatively more accurate reconstruction, with a near-rigid transformation, which may be clinically sufficient [15]. Importantly, there are no restrictions on the internal methodology: participants may choose to internally estimate a rigid transformation matrix and convert it into the four required displacement vector sets for submission.
Justification for local and global reconstruction error metrics. Local and global reconstruction errors capture complementary aspects of algorithm performance. Global reconstruction (relative to the first frame) can reveal accumulated drift over time, while local reconstruction (relative to the immediately previous frame) assesses frame-level reconstruction. These metrics are therefore indicative of both short- and long-term accuracy. Although other monotonic metrics such as final drift and Dice overlap are commonly used [7], they are excluded here to streamline evaluation. In practice, one might choose to reconstruct a sequence of ultrasound frames (as opposed to the entire scan or two adjacent frames, which are represented by local and global errors, respectively), using a pre-optimised sequence length that is most suitable to the downstream application. Since this Challenge is designed without targeting a specific clinical use case, both local and global reconstruction errors are included to span the spectrum of reconstruction performance and provide a comprehensive assessment of algorithmic accuracy.
@@ 4.4 Ranking Scheme
where the superscript denotes the normalised reconstruction error, and the subscript min and max denote the minimum and maximum errors among all participats submissions, for each corresponding metric. For each scan, the final score is computed as a weighted average of the four normalised metrics:
@@
All evaluation metrics are normalised to a common scale to prevent metrics with inherently larger magnitudes from disproportionately influencing the overall score. The two levels of measurement (global and local) and the two types of displacement vectors (pixel-based and landmark-based) are considered equally important in achieving desirable reconstruction performance. Accordingly, equal weighting is applied to each metric to establish a fair and balanced benchmark for the Challenge. A minimum score of 0 was assigned to any case where the submitted code failed to execute or the evaluation metrics cannot be computed successfully. In the event of tied overall scores, ranking was determined based on runtime. A smaller runtime was awarded a higher rank. To encourage usability in the clinical applications, a maximum runtime limit of 2 minutes per scan was enforced for all Challenge submissions. Additionally, the raw (unnormalised) values of all defined evaluation metrics were made publicly available for transparency and further analysis.
@@ 4.5 Validation and Submission
A small validation set was provided to allow participants to tune their models on previously unseen data. An example Docker template [^7] https://github.com/QiLi111/tus-rec-challenge_baseline/tree/main/submission for evaluation on the validation dataset was provided, along with the corresponding evaluation metrics. This serves to enhance the validity of the submitted Docker images and improve overall transparency in the evaluation process. Docker images from participating teams were submitted via an online form, which included a brief method description and step-by-step instructions for downloading and executing the Docker image. All submitted methods must operate in a fully automatic manner. Participants are permitted to modify and resubmit their Docker image if it fails to run on the test set due to issues such as incorrect input / output formatting or data format incompatibility. Each team was allowed a maximum of five submissions, provided the submissions represent substantively different approaches rather than minor variations in hyperparameters. The best-performing result among these were considered as the team’s final result. All submitted Docker images were independently tested by two members of the Challenge organisition team using the hidden test dataset. Evaluations were conducted on two separate platforms with identical hardware configurations: Ubuntu 18.04.6 LTS, Intel(R) Xeon(R) Gold 5215 CPU @ 2.50GHz (20 cores), NVIDIA Quadro GV100 GPU (32GB VRAM), and 128GB RAM.

### v2

@@ 4 Challenge Design
The TUS-REC2024 Challenge [^2] https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/ [^3] https://doi.org/10.5281/zenodo.10991500 is designed following the BIAS [81] Reporting Guideline for enhanced quality and transparency of biomedical research. This Challenge is associated with 5th International Workshop of Advances in Simplifying Medical UltraSound (ASMUS) at MICCAI 2024. The training and validation datasets are publicly available under CC BY-NC-SA license. The Challenge is an open-ended challenge, and submissions are welcome even after the official deadline. The test set remains held out to ensure fair benchmarking of reconstruction performance.
@@ 4.1 Task Description
Aiming at estimating the location for each ultrasound frame in 3D space, this Challenge is tasked to predict two different sets of transformation-representing dense displacement field (DDF), at global and local levels, respectively. The global DDFs denote the displacement between the current frame and the first frame, and the local DDFs represent the displacement between the current frame and the previous frame. There are no restrictions on the internal design of the algorithm, for example, whether it is learning-based, whether it processes data at the frame, sequence, or scan level, and whether it assumes rigid, affine, or non-rigid transformations.
Participating teams are provided with ultrasound sequence and corresponding transformations. Each team’s model should take an ultrasound scan as input and output four sets of displacement vectors, representing the transformations to a reference frame (i.e., the first frame or the previous frame in the sequence). During evaluation, the submitted dockerised models will be used to generate these displacement fields, from which accuracy scores will be computed to assess reconstruction performance at both local and global levels.
@@ 4.2.1 Data Collection
The dataset [^4] https://doi.org/10.5281/zenodo.11178508555https://doi.org/10.5281/zenodo.11180794666https://doi.org/10.5281/zenodo.11355499777https://doi.org/10.5281/zenodo.12752245 used in this Challenge was collected from both the left and right forearms of 85 volunteers at University College London (UCL), United Kingdom. This study was performed in accordance with the ethical standards in the 1964 Declaration of Helsinki and its later amendments or comparable ethical standards. Approval was granted by the Ethics Committee of local institution on 20th Jan. 2023 [24055/001]. The subject cohort was diverse in terms of race, gender, and age. Fig. 3 illustrates the equipment setup used during data acquisition. There were no specific exclusion criteria, except for individuals with allergies or other skin conditions that could be aggravated by the ultrasound gel. All scanned forearms were confirmed to be in healthy condition.
@@
2D ultrasound images were acquired using an Ultrasonix machine (BK, Europe) equipped with a curvilinear probe (4DC7-3/40). The ultrasound frames were captured at a rate of $20$ frames per second, with a dimension of $480\times 640$ pixels and a pixel size of around 0.2 $mm$, without applying speckle reduction as the speckle pattern may be useful for the task. Imaging was performed at a frequency of $6$ MHz, with a dynamic range of $83$ dB, an overall gain of $48\%$, and a depth of $9$ cm. Both the left and right forearms of volunteers were scanned. For each forearm, the ultrasound probe was moved along three trajectories, Straight line shape, C shape, and S shape, in both distal-to-proximal and proximal-to-distal directions. These scans were performed with the ultrasound image plane perpendicular of or parallel to the scanning direction.
In this study, an optical tracking system was chosen due to its greater accuracy and operational convenience compared to EM tracking systems. Specifically, the NDI Polaris Vicra (Northern Digital Inc., Canada) was used. During acquisition of ultrasound images, the position data recorded by the optical tracker was captured by the PLUS Toolkit [19], represented as a homogeneous transformation matrix from the tracker tool to the camera coordinate system.
As described previously in Section 2.2, the calibration matrix was obtained using the pinhead-based method [82], and the timestamps of the ultrasound frames and the transformations were aligned.
@@ 4.2.2 Sources of Errors
The primary source of error arises from the precision limitation of the optical tracker. All labels were obtained using an optical tracker with the 3D root-mean-square (RMS) volumetric accuracy acceptance criterion being less than or equal to 0.25 mm and the 3D RMS repeatability acceptance criterion being less than or equal to 0.20 mm. Slight forearm movements may occur during scanning, which is also expected in clinical environments where this technique would be deployed. These motion-induced errors are assumed to be random across different cases. Additional sources of error include inaccuracies in the calibration process (both spatial and temporal), pressure-induced skin deformation, as well as the intra-observer variability during ultrasound data acquisition, affecting probe positioning and image quality.
@@ 4.2.3 Data Pre-processing
For each scan, ultrasound frames with invalid transformation matrices, which were typically caused by occluded line of sight, were excluded. The remaining raw images, along with their corresponding transformation matrices, were temporally ordered and stored as key-value pairs in a .h5 file.
@@ 4.2.4 Data Split
Statistical power analysis was performed to determine the appropriate test sample size and minimise the likelihood of Type I and Type II errors in hypothesis testing. The effect size was calculated using Cohen’s D value [83], where the system error of the optical tracker ($0.25$ mm) was considered the meaningful difference between group means, and the standard deviation ($0.46$ mm) was derived from the results reported by [48]. A statistical power analysis for a pairwise t-test, assuming a significance level of $0.05$ and a statistical power of $0.9$, indicated a required test sample size of $31$. To ensure adequate power, we rounded up to $32$ samples ($768$ scans in total). This setup limits the probability of a Type II error to $10\%$ and a Type I error to $5\%$.
The dataset was randomly divided into training, validation, and test sets, comprising $50$, 33, and $32$ subjects, respectively. This corresponds to $1200$, $72$, and $768$ scans, amounting to $606597$, $34746$, and $384105$ frames. Ultrasound scans from the same subject will be assigned to the same set which avoids the information leak. Detailed information is described in Table 2. Specifically, the structure of the validation dataset is the same as that of the test set to ensure compatibility with the pre-defined folder hierarchy and naming conventions. This design allows submitted Docker images to run seamlessly on the test set.
@@
Protocol Statistics Train Validation Test
All # Subjects 50 3 32
# Scans 1200 72 768
# Frames 606597 34746 384105
Straight line shape # Subjects 50 3 32
# Scans 400 24 256
# Frames 192117 10515 119421
C shape # Subjects 50 3 32
# Scans 400 24 256
# Frames 202654 11655 128721
S shape # Subjects 50 3 32
# Scans 400 24 256
# Frames 211826 12576 135963
Parallel scanning # Subjects 50 3 32
# Scans 600 36 384
# Frames 298722 17228 188399
Perpendicular scanning # Subjects 50 3 32
# Scans 600 36 384
# Frames 307875 17518 195706
Left arm # Subjects 50 3 32
# Scans 600 36 384
# Frames 301155 17081 192118
Right arm # Subjects 50 3 32
# Scans 600 36 384
# Frames 305442 17665 191987
Distal-to-proximal scanning # Subjects 50 3 32
# Scans 600 36 384
# Frames 298803 16908 181844
Proximal-to-distal scanning # Subjects 50 3 32
# Scans 600 36 384
# Frames 307794 17838 202261
@@ 4.3.1 Metrics Definition
We use DDFs to evaluate the reconstruction performance, borrowing the widely recognised term used in non-rigid image registration for clarity and intuition. For each scan, participating methods are required to generate two types of DDFs representing frame-to-frame transformations, hereinafter referred to as predictions, at both global and local levels: 1) global displacement vectors are used to reconstruct all frames (excluding the first) relative to the first frame of the scan, which serves as the global reference frame; 2) local displacement vectors are used to reconstruct each frame (excluding the first) relative to its immediately previous frame, which serves as the local reference frame.
The performance of each submitted method will be assessed for every scan using two metrics: landmark reconstruction error and pixel reconstruction error: 1) landmark reconstruction error is defined as the average Euclidean distance between the ground-truth-reconstructed frame and the prediction-reconstructed frame, computed over a predefined set of landmarks; 2) pixel reconstruction error is similarly defined as the average Euclidean distance between the ground-truth and predicted reconstructions, calculated over all pixels in every frame except the first. The scale-invariant feature transform (SIFT) [84] algorithm was applied to detect landmarks. For each scan, 20 landmarks with the highest response values were selected.
Accordingly, each method should produce the following four sets of vectors:
@@
Runtime will be included as an additional evaluation metric. It is defined as the consumed time of predicting the positions for all frames but the first frame in a scan, averaged across all scans in the test set.
@@ 4.3.2 Rationale of Evaluation Metrics
Use of Euclidean distance-based error metrics vs. transformation parameter-based errors. Directly evaluating the accuracy in transformation parameter space can be biased, as the weighting of rotational and translational components can vary significantly depending on experimental setups, imaging configurations, reference coordinate systems, and definitions of rotational axes. These factors are also often application-dependent. Therefore, this Challenge adopts Euclidean distance-based metrics, which offer a less biased assessment of the discrepancy between ground truth and predicted positions in physical space.
Use of displacement-based transformation representations vs. rigid / affine matrices. Although ground-truth transformations are provided in the form of rigid transformation, we argue, based on practical experience in developing similar numerical algorithms, that requiring submissions to output homogeneous transformation matrices is not only unnecessary, but sometimes misleadingly encourages a more numerically challenging solution due to issues such as gimbal lock in using rotation matrix, local minima in numerical optimisation. In contrast, displacement-based representations allow flexibility for a quantitatively more accurate reconstruction, with a near-rigid transformation, which may be clinically sufficient [68]. Nevertheless, there are no restrictions on the internal methodology: participants may choose to internally estimate a rigid transformation matrix and convert it into the four required displacement vector sets for submission.
Justification for local and global reconstruction error metrics. Local and global reconstruction errors capture complementary aspects of algorithm performance. Global reconstruction (relative to the first frame) can reveal accumulated drift over time, while local reconstruction (relative to the immediately previous frame) assesses frame-level reconstruction. These metrics are therefore indicative of both short- and long-term accuracy. Although other monotonic metrics such as final drift and Dice overlap are also commonly used [17], they are excluded here to streamline evaluation. In practice, one might choose to reconstruct a sequence of ultrasound frames (as opposed to the entire scan or two adjacent frames, which are represented by local and global errors, respectively), using a pre-optimised sequence length that is the most suitable to specific downstream application. Since this Challenge is designed without targeting a specific clinical use case, both local and global reconstruction errors are included to span the spectrum of reconstruction performance and provide a comprehensive assessment of algorithmic accuracy.
@@ 4.4 Ranking Scheme
where the superscript denotes the normalised reconstruction error, and the subscript min{min} and max{max} denote the minimum and maximum errors among all participating submissions for each corresponding metric. For each scan, the final score is computed as a weighted average of the four normalised metrics:
@@
In this Challenge, equal weighting is used, with motivation explained later.
@@
All evaluation metrics are normalised to a common scale to prevent metrics with inherently larger magnitudes from disproportionately influencing the overall score. The two levels of measurement (global and local) and the two types of displacement vectors (pixel-based and landmark-based) are considered equally important in achieving desirable reconstruction performance. Consequently, equal weighting is applied to each metric to establish a fair and balanced benchmark for the Challenge. A minimum score of 0 was assigned to any case where the submitted code failed to execute or the evaluation metrics could not be computed successfully. In the event of tied overall scores, ranking was determined based on runtime. A smaller runtime was awarded a higher rank. To encourage usability in the clinical applications, a maximum runtime limit of 2 minutes per scan was enforced for all Challenge submissions. Additionally, the raw (unnormalised) values of all defined evaluation metrics were made publicly available for transparency and further analysis.
@@ 4.5 Validation and Submission
A small validation set was provided mainly for sanity checking on previously unseen data. An example Docker template [^8] https://github.com/QiLi111/tus-rec-challenge_baseline/tree/main/submission for evaluation on the validation dataset was provided, along with implementation of the corresponding evaluation metrics. This facilitates the preparation of valid Docker images by participants and improves overall transparency in the evaluation process. Docker images from participating teams were submitted via an online form, which included a brief method description and step-by-step instructions for downloading and executing the Docker image. All submitted methods must operate in a fully automatic manner. Participants are permitted to modify and resubmit their Docker image if it fails to run on the test set due to issues such as incorrect input/output formatting or mismatched file types and data structures. Each team was allowed to make up to five submissions, provided that each submission represents substantively different approaches rather than minor variations in hyperparameters of other submissions. The best-performing result in terms of final scores among these was considered as the team’s final result. All submitted Docker images were independently tested by two members of the Challenge organisition team using the hidden test dataset. Evaluations were conducted on two separate platforms with identical hardware configurations: Ubuntu 18.04.6 LTS, Intel(R) Xeon(R) Gold 5215 CPU @ 2.50GHz (10 cores), NVIDIA Quadro GV100 GPU (32GB VRAM), and 128GB RAM.